Let R ⊆ S × S be an equivalence relation on a set S. For an element x ∈ S, let

S(x) = {y ∈ S : (x, y) ∈ R}. Show that for every pair of elements x, y ∈ S, either

S(x) = S(y) or S(x) ∩ S(y) = ∅.



1
Expert's answer
2021-11-15T05:38:12-0500

Let "R \u2286 S \u00d7 S" be an equivalence relation on a set "S." For an element "x \u2208 S," let "S(x) = \{y ∈ S : (x, y) ∈ R\}." Let us show that for every pair of elements "x, y \u2208 S," either "S(x) = S(y)" or "S(x) \u2229 S(y) = \u2205."

Let "S(x) ∩ S(y) \ne ∅." Then there exists "z\in S(x) ∩ S(y)." Let us show that in this case "S(x) = S(y)."

Let "a\in S(x)." Then "(x,a)\in R." Since "z\in S(x)," we have that "(x,z)\in R." Taking into account that the relation "R" is symmetric, we conclude that "(z,x)\in R." Then the transitivity of "R" implies "(z,a)\in R." Since "z\in S(y)," we get that "(y,z)\in R," and transitivity of "R" implies "(y,a)\in R." Consequently, "a\in S(y)," and hence "S(x)\subseteq S(y)."

Further, let "a\in S(y)." Then "(y,a)\in R." Since "z\in S(y)," we have that "(y,z)\in R." Taking into account that the relation is symmetric, we conclude that "(z,y)\in R." Then the transitivity of "R" implies "(z,a)\in R." Since "z\in S(x)," we get that "(x,z)\in R," and transitivity of "R" implies "(x,a)\in R." Consequently, "a\in S(x)," and hence "S(y)\subseteq S(x)."

Therefore, we get that if "S(x) ∩ S(y) \ne ∅," then "S(x) = S(y)."

We conclude that for every pair of elements "x, y \u2208 S," either "S(x) = S(y)" or "S(x) \u2229 S(y) = \u2205."